# Adaptive Batch Size for Privately Finding Second-Order Stationary Points

### v1

@@ Abstract
There is a gap between finding a first-order stationary point (FOSP) and a second-order stationary point (SOSP) under differential privacy constraints, and it remains unclear whether privately finding an SOSP is more challenging than finding an FOSP. Specifically, Ganesh et al. (2023) demonstrated that an $\alpha$-SOSP can be found with $\alpha=\tilde{O}(\frac{1}{n^{1/3}}+(\frac{\sqrt{d}}{n\varepsilon})^{3/7})$, where $n$ is the dataset size, $d$ is the dimension, and $\varepsilon$ is the differential privacy parameter. Building on the SpiderBoost algorithm framework, we propose a new approach that uses adaptive batch sizes and incorporates the binary tree mechanism. Our method improves the results for privately finding an SOSP, achieving $\alpha=\tilde{O}(\frac{1}{n^{1/3}}+(\frac{\sqrt{d}}{n\varepsilon})^{1/2})$. This improved bound matches the state-of-the-art for finding an FOSP, suggesting that privately finding an SOSP may be achievable at no additional cost.
@@ 1 Introduction
Non-convex optimization has been extensively studied in recent years due to its central role in modern machine learning, and we now have a solid understanding of the complexity involved in finding FOSPs and SOSPs (Ghadimi and Lan, 2013; Agarwal et al., 2017; Carmon et al., 2020; Zhang et al., 2020). Variance reduction techniques have been shown to improve the theoretical complexity, leading to the development of several promising algorithms such as Spider (Fang et al., 2018), SARAH (Nguyen et al., 2017), and SpiderBoost (Wang et al., 2019c). More recently, private non-convex optimization has emerged as an active area of research (Wang et al., 2019b; Arora et al., 2023; Gao and Wright, 2023; Ganesh et al., 2023; Lowy et al., 2024; Menart et al., 2024).
@@ 1.1 Our Main Result
Given the dataset size of $n$, privacy parameters $\varepsilon,\delta$, and functions defined over $d$-dimension space, Ganesh et al. (2023) proposed a private algorithm that can find an $\alpha_{S}$-SOSP for $F_{\mathcal{P}}$ with
@@
When the privacy parameter $\varepsilon$ is sufficiently small, we observe that $\alpha_{F}\ll\alpha_{S}$. This raises the question: is finding an SOSP under differential privacy constraints more difficult than finding an FOSP?
This work improves the results of Ganesh et al. (2023). Specifically, we present an algorithm that finds an $\alpha$-SOSP with privacy guarantees, where:
@@ 1.3 Other Related Work
A significant body of literature on private optimization focuses on the convex setting, where it is typically assumed that each function $f(;z)$ is convex for any $z$ in the universe (e.g., (Chaudhuri et al., 2011; Bassily et al., 2014, 2019; Feldman et al., 2020; Asi et al., 2021; Kulkarni et al., 2021; Carmon et al., 2023)).
The tree mechanism, originally introduced by the differential privacy (DP) community (Dwork et al., 2010; Chan et al., 2011) for the continual observation, has inspired tree-structure private optimization algorithms like Asi et al. (2021); Bassily et al. (2021); Arora et al. (2023); Zhang et al. (2024). Some prior works have explored adaptive batch size techniques in optimization. For instance, De et al. (2016) introduced adaptive batch sizing for stochastic gradient descent (SGD), while Ji et al. (2020) combined adaptive batch sizing with variance reduction techniques to modify SVRG and Spider algorithms. However, these works’ motivations and approaches to setting adaptive batch sizes differ from ours. To the best of our knowledge, we are the first to propose using adaptive batch sizes in the context of optimization under differential privacy constraints.
@@ 2.1 Other Techniques
Let $\mathcal{Z}_{1},\cdots,\mathcal{Z}_{\Sigma}$ be dataset spaces, and $\mathcal{X}$ be the state space. Let $\mathcal{M}_{i}:\mathcal{X}^{i-1}\times\mathcal{Z}_{i}\to\mathcal{X}$ be a sequence of algorithms for $i\in[\Sigma]$. Let $\mathcal{A}:\mathcal{Z}^{(1:\Sigma)}\to\mathcal{X}^{\Sigma}$ be the algorithm that given a dataset $Z_{1:\Sigma}\in\mathcal{Z}^{(1:\Sigma)}$, sequentially computes $X_{i}=\sum_{j=1}^{i}\mathcal{M}_{i}(X_{1:j-1},Z_{i})+\mathrm{TREE}(i)$ for $i\in[\Sigma]$, and then outputs $X_{1:\Sigma}$.
@@
Definition 2.3 (SubGaussian, and Norm-SubGaussian).
Theorem 2.5 (Matrix Bernstein Inequality, Tropp et al. (2015)).
Consider a finite sequence of independent, random matrices $X_{1},\cdots,X_{n}$ with common dimensions $d\times d$ and $\operatorname*{\mathbb{E}}[X_{i}]=0,\|X_{i}\|\leq L$ a.s., $\forall i$. Let $\sigma^{2}=\|\sum_{i=1}^{n}\operatorname*{\mathbb{E}}[X_{i}^{2}]\|$. Let $S=\sum_{i=1}^{n}X_{i}$. Then for any $t\geq 0$, we have
$\Pr[\|S\|\geq t]\leq d\cdot\exp(-\frac{-t^{2}}{\sigma^{2}+Lt/3}).$
1:Input: Noise parameter $\sigma$, sequence length $\Sigma$
@@ 3 SOSP
Assume $F$ is $M$-smooth and let $\eta=1/M$. Let $x_{t+1}=x_{t}-\eta\widetilde{\nabla}$. Then we have $F(x_{t+1})\leq F(x_{t})+\eta\|\widetilde{\nabla}_{t}\|\cdot\|\nabla F(x_{t})-\widetilde{\nabla}_{t}\|-\frac{\eta}{2}\|\widetilde{\nabla}_{t}\|^{2}.$ Moreover, if $\|\nabla F(x_{t})\|\geq\gamma$ and ~$\|\widetilde{\nabla}_{t}-\nabla F(x_{t})\|\leq\gamma/4$, we have
$F(x_{t+1})-F(x_{t})\leq-\eta\|\widetilde{\nabla}_{t}\|^{2}/16.$
@@
$=F(x_{t})-\eta/2\|\widetilde{\nabla}_{t}\|^{2}-\langle\nabla F(x_{t})-\widetilde{\nabla}_{t},\eta\widetilde{\nabla}_{t}\rangle$
$\leq F(x_{t})+\eta\|\nabla F(x_{t})-\widetilde{\nabla}_{t}\|\cdot\|\widetilde{\nabla}_{t}\|-\frac{\eta}{2}\|\widetilde{\nabla}_{t}\|^{2}.$
When $\|\nabla F(x_{t})\|\geq\gamma$ and ~$\|\widetilde{\nabla}_{t}-\nabla F(x_{t})\|\leq\gamma/4$, the conclusion follows from the calculation.
One of the main challenges to finding the SOSP, compared with finding the FOSP, is showing the algorithm can escape from saddle points, which was addressed by previous results (e.g.,Jin et al. (2017)) and was established in the DP literature as well:
Lemma 3.3 (Essentially from Wang et al. (2019a)).
Under Assumption 3.1, run SGD iterations $x_{t+1}=x_{t}-\eta\nabla_{t}$, with step size $\eta=1/M$. Suppose $x_{0}$ is a saddle point satisfying $\|\nabla F(x_{0})\|\leq\alpha$ and $\textup{smin}(\nabla^{2}F(x_{0}))\leq-\sqrt{\rho\alpha}$, $\alpha=\gamma\log^{3}(dBM/\rho\iota)$. If $\nabla_{0}=\nabla F(x_{0})+\zeta_{1}+\zeta_{2}$ where $\|\zeta_{1}\|\leq\gamma$, $\zeta_{2}\sim\mathcal{N}(0,\frac{\gamma^{2}}{d\log(d/\iota)}I_{d})$, and $\|\nabla_{t}-\nabla F(x_{t})\|\leq\gamma$ for all $t\in[\Gamma]$, with probability at least $1-\iota$, one has $F(x_{\Gamma})-F(x_{0})\leq-\Omega\big{(}\frac{\gamma^{3/2}}{\sqrt{\rho}\log^{3}(\frac{dMB}{\rho\gamma\iota})}\big{)},$ where $\Gamma=\frac{M\log(\frac{dMB}{\rho\gamma\iota})}{\sqrt{\rho\gamma}}$.
@@
Lemma 3.3 suggests that, if we meet a saddle point, then after the following $\tilde{O}(1/\sqrt{\gamma})$ steps, the function value will decrease by at least $\tilde{\Omega}(\gamma^{3/2})$. This means the function value decreases by $\tilde{\Omega}(\gamma^{2})$ on average for each step. Similar to the proof in Ganesh et al. (2023), we have the following guarantee of Stochastic SpiderBoost:

### v2

@@ Abstract
There is a gap between finding a first-order stationary point (FOSP) and a second-order stationary point (SOSP) under differential privacy constraints, and it remains unclear whether privately finding an SOSP is more challenging than finding an FOSP. Specifically, Ganesh et al. (2023) claimed that an $\alpha$-SOSP can be found with $\alpha=\tilde{O}(\frac{1}{n^{1/3}}+(\frac{\sqrt{d}}{n\varepsilon})^{3/7})$, where $n$ is the dataset size, $d$ is the dimension, and $\varepsilon$ is the differential privacy parameter. However, a recent analysis revealed an issue in their saddle point escape procedure, leading to weaker guarantees. Building on the SpiderBoost algorithm framework, we propose a new approach that uses adaptive batch sizes and incorporates the binary tree mechanism. Our method not only corrects this issue but also improves the results for privately finding an SOSP, achieving $\alpha=\tilde{O}(\frac{1}{n^{1/3}}+(\frac{\sqrt{d}}{n\varepsilon})^{1/2})$. This improved bound matches the state-of-the-art for finding a FOSP, suggesting that privately finding an SOSP may be achievable at no additional cost.
@@ 1 Introduction
Non-convex optimization has been extensively studied in recent years due to its central role in modern machine learning, and we now have a solid understanding of the complexity involved in finding FOSPs and SOSPs (Ghadimi and Lan, 2013; Agarwal et al., 2017; Carmon et al., 2020; Zhang et al., 2020). Variance reduction techniques have been shown to improve the theoretical complexity, leading to the development of several promising algorithms such as Spider (Fang et al., 2018), SARAH (Nguyen et al., 2017), and SpiderBoost (Wang et al., 2019b). More recently, private non-convex optimization has emerged as an active area of research (Wang et al., 2019a; Tran and Cutkosky, 2022; Arora et al., 2023; Gao and Wright, 2023; Ganesh et al., 2023; Wang et al., 2023; Lowy et al., 2024; Kornowski et al., 2024; Menart et al., 2024).
@@ 1.1 Our Main Result
Given the dataset size of $n$, privacy parameters $\varepsilon,\delta$, and functions defined over $d$-dimensional space, Ganesh et al. (2023) proposed a private algorithm that can find an $\alpha_{S}$-SOSP for $F_{\mathcal{P}}$ with
@@
When the privacy parameter $\varepsilon$ is sufficiently small, and the error term depending on it dominates the non-private term $1/n^{1/3}$, we observe that $\alpha_{F}\ll\alpha_{S}$. This raises the question: is finding an SOSP under differential privacy constraints more difficult than finding an FOSP?
Moreover, as pointed out by Tao et al. (2025), the results in Ganesh et al. (2023) may be overly optimistic. In particular, the saddle point escape subprocedure used in Ganesh et al. (2023) was designed for the full-batch setting. However, to mitigate dependence issues, the algorithm performs only a single pass over the $n$ functions and relies on minibatches instead. A direct fix for this issue leads to a weaker guarantee on the minimum eigenvalue of the Hessian, specifically,
$\nabla^{2}F_{\mathcal{P}}(x)\succeq-\sqrt{\rho}d^{1/5}\alpha^{2/5}I_{d},$
which fails to satisfy Definition 1.2. To address this, Tao et al. (2025) proposed an alternative correction, but their method resulted in weaker theoretical guarantees than Ganesh et al. (2023) originally claimed, yielding
$\alpha=\tilde{O}\big{(}\frac{1}{n^{1/3}}+(\frac{\sqrt{d}}{n\varepsilon})^{2/5}\big{)}.$
This work improves upon the results of Ganesh et al. (2023). In addition, we introduce a new saddle point escape subprocedure that correctly addresses the issue in Ganesh et al. (2023) while fully recovering—and even strengthening—the theoretical guarantees.
Specifically, we present an algorithm that finds an $\alpha$-SOSP with privacy guarantees, where:
@@ 1.2 Overview of Techniques
Fixing the Error:
Building upon our discussion of obtaining more accurate gradient estimations through adaptive batch sizes, we extend this approach to Hessian estimations, achieving accuracy in terms of the operator norm. Upon encountering a potential saddle point—characterized by a small gradient norm—we utilize the Hessian to inform the gradient estimation over several iterations. This process is analogous to performing stochastic power iteration methods on the Hessian to identify the direction corresponding to the smallest eigenvalue. If the Hessian exhibits a small enough eigenvalue (say smaller than $-\sqrt{\rho\alpha}$), we demonstrate that this approach facilitates a significant drift, effectively enabling successful escape from the saddle point. This idea can also be applied to Ganesh et al. (2023) to recover their claimed rate.
@@ 1.3 Other Related Work
A significant body of literature on private optimization focuses on the convex setting, where it is typically assumed that each function $f(;z)$ is convex for any $z$ in the universe (e.g., (Chaudhuri et al., 2011; Bassily et al., 2014, 2019; Feldman et al., 2020; Asi et al., 2021; Kulkarni et al., 2021; Carmon et al., 2023; Gopi et al., 2023)).
The tree mechanism, originally introduced by the differential privacy (DP) community (Dwork et al., 2010; Chan et al., 2011) for the continual observation, has inspired tree-structure private optimization algorithms like Asi et al. (2021); Bassily et al. (2021); Arora et al. (2023); Zhang et al. (2024). One can also use the matrix mechanism Fichtenberger et al. (2023) which improves the tree mechanism by a constant factor. Some prior works have explored adaptive batch size techniques in optimization. For instance, De et al. (2016) introduced adaptive batch sizing for stochastic gradient descent (SGD), while Ji et al. (2020) combined adaptive batch sizing with variance reduction techniques to modify SVRG and Spider algorithms. However, these works’ motivations and approaches to setting adaptive batch sizes differ from ours. To the best of our knowledge, we are the first to propose using adaptive batch sizes in the context of optimization under differential privacy constraints.
@@ 2.1 Other Techniques
Let $\mathcal{Z}_{1},\cdots,\mathcal{Z}_{\Sigma}$ be dataset spaces, and $\mathcal{X}$ be the state space. Let $\mathcal{M}_{i}:\mathcal{X}^{i-1}\times\mathcal{Z}_{i}\to\mathcal{X}$ be a sequence of algorithms for $i\in[\Sigma]$. Let $\mathcal{A}:\mathcal{Z}^{(1:\Sigma)}\to\mathcal{X}^{\Sigma}$ be the algorithm that given a dataset $Z_{1:\Sigma}\in\mathcal{Z}^{(1:\Sigma)}$, sequentially computes $X_{i}=\sum_{j=1}^{i}\mathcal{M}_{j}(X_{1:j-1},Z_{j})+\mathrm{TREE}(i)$ for $i\in[\Sigma]$, and then outputs $X_{1:\Sigma}$.
@@
1:Input: Noise parameter $\sigma_{\mathrm{tree}}$, sequence length $\Sigma$
@@
Definition 2.3 (SubGaussian, and Norm-SubGaussian).
Theorem 2.5 (Matrix Azuma Inequality, Tropp (2012)).
Consider a finite sequence of random self-adjoint matrices $X_{1},\cdots,X_{n}$ with common dimensions $d\times d$ and $\operatorname*{\mathbb{E}}[X_{i}\mid\mathcal{F}_{i-1}]=0,X_{i}\preceq A_{i}$ a.s., $\forall i$. Let $\sigma^{2}=\|\sum_{i=1}^{n}A_{i}^{2}\|$. Let $S=\sum_{i=1}^{n}X_{i}$. Then for any $t\geq 0$, we have
$\Pr[\|S\|\geq t]\leq d\cdot\exp(-\frac{-t^{2}}{8\sigma^{2}}).$
@@ 3 SOSP
Assume $F$ is $M$-smooth and let $\eta=1/M$. Let $x_{t+1}=x_{t}-\eta g_{t}$. Then we have $F(x_{t+1})\leq F(x_{t})+\eta\|g_{t}\|\cdot\|\nabla F(x_{t})-g_{t}\|-\frac{\eta}{2}\|g_{t}\|^{2}.$ Moreover, if $\|\nabla F(x_{t})\|\geq\gamma$ and $\|g_{t}-\nabla F(x_{t})\|\leq\gamma/4$, we have
$F(x_{t+1})-F(x_{t})\leq-\eta\|g_{t}\|^{2}/16.$
@@
$=F(x_{t})-\eta/2\|g_{t}\|^{2}-\langle\nabla F(x_{t})-g_{t},\eta g_{t}\rangle$
$\leq F(x_{t})+\eta\|\nabla F(x_{t})-g_{t}\|\cdot\|g_{t}\|-\frac{\eta}{2}\|g_{t}\|^{2}.$
When $\|\nabla F(x_{t})\|\geq\gamma$ and $\|g_{t}-\nabla F(x_{t})\|\leq\gamma/4$, the conclusion follows from the calculation. ∎
Lemma 3.2 shows that one can be able to find the FOSP with inexact gradient estimates as long as the estimated error is small enough. A key challenge in finding an SOSP, compared to an FOSP, is ensuring that the algorithm can effectively escape saddle points. Existing analyses of saddle point escape using inexact gradients are insufficient for our purposes. To address this, we propose a new approach that leverages both inexact gradients and Hessians for escaping saddle points. The details are presented in the following subsection.
3.1 Escape Saddle Point with Hessian
We present the subprocedure for escaping saddle points in this section, with its pseudocode provided in Algorithm 2. The core idea is straightforward: given a sufficiently accurate estimate of the function’s Hessian, we can apply the power method to escape the saddle point. If the Hessian’s smallest eigenvalue is sufficiently negative, then after a certain number of steps, the iterate will have moved a significant distance, resulting in a substantial decrease in function value—indicating a successful escape.
Definition 3.3.
We say the estimation $(g,H)$ of the pair of gradient and Hessian is $(\gamma,\varkappa)$-accurate at point $x$, if
$\|g-\nabla F(x)\|_{2}\leq\gamma,\|H-\nabla^{2}F(x)\|_{2}\leq\varkappa.$
Algorithm 2 Escape from Saddle Point
1:Input: initial point $x_{0}$ such that $\|\nabla F(x)\|\leq\alpha$, $(\gamma,\varkappa)$-accurate estimation pair $(g,H)$ at $x_{0}$, parameters $\Gamma,\Xi$
2:Process:
3:$y\leftarrow x_{0}$
4:for $t=1,\cdots,\Gamma$ do
5:     $g_{t-1}=g+H(x_{t-1}-x_{0}))+\zeta_{t-1}$, where $\zeta_{t-1}\sim\mathcal{N}(0,\sigma_{t-1}^{2}I_{d})$.
6:     $x_{t}=x_{t-1}-\eta g_{t-1}$
7:     if $\|x_{t}-x_{0}\|\geq\Xi$ then
8:         $y\leftarrow x_{t}$
9:         Break
10:     end if
11:end for
12:Output: $y$
The Gaussian noise $\zeta_{t-1}$ added in Line 5 is for the privacy purpose. We have the following guarantee of Algorithm 2:
Theorem 3.4.
Suppose the function $F$ satisfies the Assumption 3.1, and the initial point is a saddle point such that $\|\nabla F(x)\|\leq\alpha$ and $\nabla^{2}F(x)\succeq-\sqrt{\rho\alpha}I_{d}$. When $\alpha\geq\gamma\log^{3}(dBM/\rho\iota)$, $\sigma_{t}=\frac{\gamma}{\sqrt{d\log(T/\iota)}}$ $\Gamma=\tilde{O}(M/\sqrt{\rho\gamma})$, $\varkappa\Xi\leq\gamma,\eta=1/M$ and $\Xi=\sqrt{\gamma/\rho}$, then with probability at least $1-\iota$, the output $y$ satisfies that
$F(y)-F(x_{0})\leq-\Phi,$
where $\Phi=\Omega(\frac{\gamma^{3/2}}{\sqrt{\rho}\log^{3}(dMZb/\rho\gamma\iota)})$.
We have the following guarantee on $g_{t}$:
Lemma 3.5.
With probability at least $21-\iota/$ 2, for all $t\in[\Gamma]$, we have
$\|g_{t-1}-\nabla F(x_{t-1})\|_{2}\leq 2\gamma+(\varkappa+\rho\Xi)\Xi.$
Let $\varrho_{t}=\nabla F(x_{t})-g_{t}$ denote the difference between the true gradient and our estimation. We use the following standard technical claims to connect the distance of the trajectory and the function value change.
Claim 3.6 (Lemma 10 in Wang et al. (2019a)).
We have
$F(x_{t+1})-F(x_{t})\leq-\frac{\eta}{4}\|\nabla F(x_{t})\|_{2}^{2}+5\eta\|\varrho_{t}\|_{2}^{2},$
Claim 3.7 (Lemma 11 in Wang et al. (2019a)).
For any $t\in[\Gamma]$, one has
$\|x_{t}-x_{0}\|_{2}^{2}\leq 8\eta\Gamma(F(x_{0})-F(x_{t}))+50\eta^{2}t\sum_{i=1}^{t}\|\varrho_{i}\|_{2}^{2}.$
In Algorithm 2, we halt the algorithm whenever we find a point $x_{t}$ such that $\|x_{t}-x_{0}\|_{2}\geq\Xi$. We use the following lemma to demonstrate that a large distance at $x_{t}$ means a large function value decrease at $x_{t}$, that is, it escapes from the saddle point successfully:
Lemma 3.8.
Suppose Algorithm 2 halts in advance when condition $\|x_{t}-x_{0}\|\geq\Xi$ is satisfied, then we have
$F(x_{t})-F(x_{0})\leq-\Phi.$
Similar to previous works Jin et al. (2017); Wang et al. (2019a), we use a coupling argument to prove that we can escape the saddle point with high probability. Fix the sequence of Gaussian noise $(\zeta_{1},\cdots,\zeta_{\Gamma})$ and ensure $\|\zeta_{t}\|_{2}\leq\gamma$ for all $t\in[\Gamma]$.
Let $y(x)$ denote the output $y$ conditional on the first iterate $x_{1}=x$. Define the set of stuck region around $x_{0}-\eta g$:
$\mathcal{X}(x_{0})=\{x\mid x\in B(x_{0}-\eta g,r),\text{ and }\|y(x)-x_{0}\|_{2}\leq\Xi\},$ (1)
where $r=O(\eta\gamma)$. Suppose $x_{0}$ is the saddle point, and let $\mathbf{e}_{1}$ be the minimum eigenvector of $H$. We have the following Lemma:
Lemma 3.9.
Suppose $\varkappa\leq\sqrt{\rho\gamma}/2$. For any two points $w,u\in B(x_{0},r)$, if $w-u=\mu r\mathbf{e}_{1}$ with $\mu\geq\iota^{2}/16\sqrt{d}$, then at least one of $w,u$ is not in $\mathcal{X}(x_{0})$.
@@ 3.1 Escape Saddle Point with Hessian
With these Lemmas, we can complete the proof of Theorem 3.4. In particular, Theorem 3.4 suggests that, if we meet a saddle point, then after the following $\tilde{O}(1/\sqrt{\gamma})$ steps, the function value will decrease by at least $\tilde{\Omega}(\gamma^{3/2})$. This means the function value decreases by $\tilde{\Omega}(\gamma^{2})$ on average for each step.
3.2 Main Algorithm
Algorithm 3 Stochastic Spider with Escaping Saddle Point SubProcedure
1:Input: Dataset $\mathcal{D}$, privacy parameters $\varepsilon,\delta$, parameters of objective function $B,M,G,\rho$, parameter $\kappa$, failure probability $\omega$, batch size parameter $b$, noise parameters $\sigma_{\mathrm{tree}},\{\sigma_{t}\}$
2:set $\mathrm{drift}_{0}=\kappa,\mathrm{frozen}_{-1}=1,\Delta_{-1}=0,\mathcal{D}_{r}\leftarrow\mathcal{D},t=0,\mathrm{EscapeFlag}=\mathbf{False}$
3:while $t<T$ and the number of unused functions is larger than $b$ do
4:   // Estimate gradient and Hessian with oracle (Algorithm 4)
5:     if $\mathrm{drift}_{t}\geq\kappa$ or $\mathrm{count}\geq\tau$ then
6:         $\nabla_{t}=\mathcal{O}_{1}(x_{t})$, $H_{t}=\mathcal{O}_{3}(x_{t})$,
7:         $\mathrm{drift}_{t}=0$, $\mathrm{frozen}_{t}=\mathrm{frozen}_{t-1}-1$, $\mathrm{count}=0$
8:     else
9:         $\Delta_{t}=\mathcal{O}_{2}(x_{t},x_{t-1})$, $\nabla_{t}=\nabla_{t-1}+\Delta_{t}$, ~$\widetilde{\nabla}_{t}=\nabla_{t}+\mathrm{TREE}(t)$
10:         $\Delta_{t}^{H}=\mathcal{O}_{4}(x_{t},x_{t-1})$, $H_{t}=H_{t-1}+\Delta_{t}^{H}$
11:     end if
12:// Escape Saddle Point if $\mathrm{EscapeFlag}=\mathbf{True}$
13:     if ~$\|\widetilde{\nabla}_{t}\|\leq\gamma\log^{3}(BMd/\rho\omega)\bigwedge\mathrm{frozen}_{t-1}\leq 0$ then
14:         Set $\mathrm{frozen}_{t}=\Gamma$, $\mathrm{EscapeFlag}=\mathbf{True}$, $g=\widetilde{\nabla}_{t}$, $H=H_{t}$, $x_{\mathrm{anchor}}=x_{t}$, $\mathrm{count}=\mathrm{count}+1$
15:     end if
16:     if $\mathrm{EscapeFlag}==\mathbf{True}$ then
17:         $g_{t}=g+H(x_{t}-x_{\mathrm{anchor}})+\xi_{t}$, where $\xi_{t}\sim\mathcal{N}(0,\sigma_{t}^{2}I_{d})$
18:     else
19:         $g_{t}=\widetilde{\nabla}_{t}$triangleright Normal gradient descent
20:     end if
21:     $x_{t+1}=x_{t}-\eta g_{t},\mathrm{drift}_{t+1}=\mathrm{drift}_{t}+\|g_{t}\|_{2}$, $\mathrm{frozen}_{t+1}=\mathrm{frozen}_{t}-1$,
22:     if $\|x_{t+1}-x_{\mathrm{anchor}}\|\geq\Xi$ or $\mathrm{frozen}_{t+1}\leq 0$ then
23:         $\mathrm{EscapeFlag}=\mathbf{False}$
24:     end if
25:     $t=t+1$
26:end while
27:Return: $\{x_{1},\cdots,x_{t}\}$
Algorithm 3 follows the SpiderBoost framework. We primarily focus on gradient oracles and estimators for simplicity, as the Hessian case follows the same principle. Moreover, we allow a larger error tolerance for a Hessian estimate $H_{t}$, approximately $\sqrt{\rho\gamma}$, compared to the gradient, which is $\gamma$. We discuss some key variables and parameters in it.
We either query $\mathcal{O}_{1}$ to estimate the gradient directly or query $\mathcal{O}_{2}$ to estimate the gradient difference between consecutive points. The term $\mathrm{drift}$ controls the estimated error: when $\mathrm{drift}_{t}$ is small, $\Delta_{t}$ remains a reliable estimator; when $\mathrm{drift}_{t}$ exceeds the threshold determined by the parameter $\kappa$, we obtain a fresh estimate from $\mathcal{O}_{1}$.
The term $\mathrm{frozen}$ serves a technical role in the application of Theorem 3.4; specifically, when a potential saddle point is detected, we invoke the subprocedure described in Algorithm 2.
The variable $\mathrm{count}$ tracks the number of saddle point escapes since the last fresh gradient estimate $\Delta_{t}$ and Hessian estimate $H_{t}$. Once $\mathrm{count}$ exceeds the threshold $\tau$, we force a fresh query to $\mathcal{O}_{1}$ and $\mathcal{O}_{3}$ to ensure privacy, following the Gaussian Mechanism.
We have the following guarantee of Algorithm 3.